IWOTA – 2000 Draft program

Tuesday, 13 June

8.30–9.30  Registration
9.40–9.45  Opening

Chairman: Kaashoek

9.45–10.35  Meyer

Coffee

Chairman: Sadosky

11.00–11.50  Curtain
11.55–12.45  Treil

Lunch

Section A Section B Section C Section D Section E
Chairmen:  Amar  Alpay  Spitkovsky Tretter  Yger
14.35–15.20  Douglas  Kheifets  Leblond  Mityagin  Staffans
15.25–16.10  Lomonosov  Ball  Dym  Naboko  Putinar

Coffee

Chairman: Volberg

16.45–17.35  De Branges
17.45  Departure to the reception

Wednesday, 14 June

Chairman: Dym

8.40–9.30  Foias
9.35–10.25  Alpay

Coffee

Chairman: H. Langer

10.50–11.40  Birman
11.45–12.35  Megretski

Lunch

Section A Section B Section C Section D Section E
Chairmen: Douglas  Seip  Ferguson  Tkachenko Pavlov
14.25–15.10  Le Merdy McCarty  Arazy  Solomyak Latushkin
15.15–15.45  Li  Kapustin Yu. Karlovich Mennicken Minkin
15.50–16.20  Abakumov Bakonyi  Dritschel  Petkov  Belishev

Coffee

Chairmen: Le Merdy McCarthy  Böttcher  Shkalikov  Putinar
16.45–17.15  Amar  Bruzual  Brudnyi  Gasimov  Jaming
17.20–17.50  Tomilov  Kerchy  Nakazi  Berezanski  Yger
17.55–18.25  Zemanek Dehghan  Lyubarskii Olshevski  Xia
18.30–19.00  Kurasov  Kosiek  Pott  Hasanov  de Snoo

Thursday, 15 June

Chairman: Partington

8.40–9.30  Böttcher
9.35–10.25  Sadosky

Coffee

Chairman: Staffans

10.50–11.40  Volberg
11.45–12.35  Saff

Lunch

Section A Section B Section C Section D Section E
Chairmen: Rodman Vasilescu  Treil  Megretski Vinnikov
14.25–15.10  Kaashoek Shimorin  Spitkovsky  Shkalikov  Pavlov
15.15–16.00  Nevanlinna Partington  Ferguson  Tkachenko Yafaev

Coffee

Chairmen:  Peller  Baratchart Verdyun Lunel Naboko  Yafaev
16.25–16.55  Azizov  Dyakonov  Litvinchuk  Malamud  Ivanov
17.00–17.30  Markus  Badea  Matsaev  Janas  Nudelman
17.35–18.05  Tikhonov  Kupin  Paneah  van Shagen Gubreev
18.15–19.40  Poster session/Coffee
20.00  Dinner

Friday, 16 June

Chairman: Helton

8.40–9.30  Verduyn Lunel
9.35–10.25  Seip

Coffee

Chairman: Mennicken

10.50–11.40  Pisier
11.45–12.35  Davidson

Lunch

Section A Section B Section C Section D Section E
Chairmen: Cassier  Saff  Davidson Mityagin  Curtain
14.25–15.10  Peller  Yuditskii  Coburn  Arov  Vinnikov
15.15–16.00  Adamyan  H. Langer Poltoratskii Zworski  Larson
Chairmen: Nevanlinna  Deville  Coburn  Petkov  Solomyak
16.05–16.35  Belitskii Tsekanovskii  Xu  Tretter  Fedorov
16.40–17.10  Rodman Marcantognini Pedersen  Zwart  Avdonin

Coffee

Chairman: Kaashoek

17.45–18.35  Helton
18.35  Closing